From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 28 2003 - 07:47:57 EST
------------------------------------------------------------ Date: Tue Jan 28 07:47:57 2003, Shift: 00:00-08:00 Supervisor: BB, with: IA Log book#13 pages 1 - 2 Shift Summary: ------------------------------------------------------------ 1. MRS was at 90 degrees, 6B; FS settings: 8 degrees, 1/4A. 2. Run 7676 containing 73k trigger 6 and 80k trigger 3 events, but the backround was high. At the begining the TofW/ZDC ratio was 17 and increased to 25 at the end of the run. Voted to dump store at 5:45 am. 3. No problem detected. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Tue Jan 28 2003 - 07:53:13 EST